Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1.
1
Expert's answer
2019-07-18T10:25:02-0400
"\theta=-\frac{3\pi}{4}"

"tan\theta = 1 =\frac{\text{y coordinate}}{\text{x coordinate}}=-\frac{\text{y coordinate}}{1}"

"\text{y coordinate}=-1"

"P(-1,-1)"